OpenStudy (anonymous):

When solving the equation 4m - 8 = -12, what is the next logical step and why?

OpenStudy (anonymous):

4m = -4; Addition Property of Equality

OpenStudy (anonymous):

correct?

OpenStudy (yacoub1993):

\[4m-8=-12\] \[4m=-12+8\] \[4m=-4\] \[m=-1\]

OpenStudy (yacoub1993):

Correct